OpenStudy (anonymous):

6x-3y=3; (1,1), (-1,-3)

OpenStudy (anonymous):

all this is, is plugging in the points so for (1,1) it's actually (x,y) right? so x=1 and y=1 now all you have to do is plug in points. 6x-3y=3 6(1)-3(1)=3 6-3=3 3=3 so this is correct. now onto the second set of points. (-1,-3) is (x,y) so x=-1 and y=-3 6x-3y=3 6(-1)-3(-3)=3 -6+6=3 ok here i'm confused, is this what you wanted us to do?

OpenStudy (anonymous):

yes and thank you!!